# Frequency-Guided Masking for Enhanced Vision Self-Supervised Learning

### v1

@@ 4.2.2 Few Shot Learning
| Method | Base LR = 2e-4 | Base LR = 2e-4 | Base LR = 2e-3 | AVG |
| Warm Up = 0 | Warm Up = 100 | Warm Up = 5 | ||
| iBOT | 64.0 | 71.1 | 2.0 | 45.7 |
@@
| MFM | 57.7 | 58.5 | 41.9 | 52.7 |
[…]
| FOLK | 71.2 | 68.1 | 62.2 | 67.2 |
Table 2 demonstrates the potential of our approach to effectively leverage small datasets, making it especially relevant for applications where data is scarce. The MFM model does not perform well with limited data in downstream tasks assumably because, during the pretext task, MFM does not see the original image without augmentation. And this limitation has been addressed in our proposed FOLK method, owing to the self-distillation design. FOLK is more robust in different training settings compared to other methods.
4.2.3 Semantic Segmentation
@@
Results of this fine-tuning effort for the semantic segmentation task on the ADE20K dataset are shown in Table 3. It presents a comparative analysis of different methodologies: Supervised Learning, iBOT, iBOT with Attention Mask (AttMask), MFM, and our FOLK model at different pre-training epochs ( and epochs). Notably, the FOLK model with epochs of pre-training achieves the highest mIoU at , slightly surpassing the iBOT’s mIoU. This indicates a successful adaptation of the FOLK methodology, showing not only an improvement over the MFM results but also demonstrating that extended pre-training can lead to marginal yet significant performance gains. Furthermore, even at pre-training epochs, FOLK performs comparably to other advanced methods, highlighting its efficacy in leveraging the dataset and architecture for semantic segmentation tasks.
4.2.4 Ablation Study
@@
In another ablation study, we examine the effects of employing various filters for frequency masking, which demonstrates the superior effectiveness of our informed filters design. This part is detailed in Appendix B.3).
5 Conclusion
Appendix
The following sections present a comprehensive overview of the FOLK implementation, detailing both the pre-training and fine-tuning phases. We also provide extensive experimental analyses, including performance evaluations on CNN-based model image classification, few-shot learning, and method efficiency analysis. Additionally, we will showcase visualizations of our designed informed filters for various input images, as well as the outputs generated by our pre-trained model on the pretext task, further demonstrating the effectiveness of our approach.
Appendix A Implementation Details
Throughout our experimental investigations, we adopted the methodologies prescribed by the iBOT [Zhou et al., 2022] and MFM [Xie et al., 2023] frameworks. In a multi-GPU training circumstance, the learning rate adjustment is vital for optimizing the model’s learning efficiency. The formulation used to compute the scaled learning rate is delineated below:
In this context, signifies the optimal or peak learning rate identified for the model’s training process. The term refers to the number of training examples processed simultaneously by each GPU. Lastly, denotes the total count of GPUs employed for parallel computation. The coefficient in the denominator is a normalization factor, ensuring the scaled learning rate maintains an appropriate magnitude relative to the hardware configuration. We used the PyTorch Library [Paszke et al., 2019] for our code development.
A.1 Pre-Train Stage
Our pre-training procedures largely align with those outlined in the BEiT [Bao et al., 2022] study, albeit with a few modifications. Specifically, our pre-training regime for the models incorporates simple yet effective data augmentation techniques, including random resized cropping to a resolution of pixels and image flipping.
We employ the AdamW optimizer [Loshchilov and Hutter, 2019], with a pre-training duration set to or epochs, a batch size of , per GPU, and a peak learning rate of . Additional parameters include a cosine decay learning rate schedule, warmup epochs, and a specific setting for optimizer momentum (, = , ) [Chen et al., 2020a] with a weight decay of . Also, we used a value of for gradient clipping to prevent the exploding gradient problem.
A.2 Fine-Tune Stage
For the fine-tuning stage, we tried to keep most of the configuration of our FOLK the same as MFM [Xie et al., 2023] for a fair comparison.
A.2.1 Classification Task
We ran epochs for fine-tuning the pre-trained model (i.e. ViT-S/16) on ImageNet-1K for image classification, employing the optimizer across all configurations with a weight decay of and the optimizer momentum , = , . Moreover, the approach includes a cosine decay learning rate schedule [Li and Arora, 2020], with a layer-wise learning rate decay equal to [Bao et al., 2022, Clark et al., 2020]. We also utilized advanced augmentation techniques such as Mixup [Zhang et al., 2018] and Cutmix [Yun et al., 2019], as well as label smoothing and random augmentation to further improve model robustness and generalization capability [Szegedy et al., 2016, Cubuk et al., 2020]. The batch size is maintained at , with a peak learning rate set at .
In contrast, the fine-tuning settings for the ResNet-50 model [He et al., 2016] generally follow the configurations suggested by [Wightman et al., 2021], with modifications to adopt the AdamW optimizer as recommended by [Fang et al., 2023]. This includes a binary cross-entropy loss function [Zhang and Sabuncu, 2018] and adjustments to the learning rate scheduler. The weight decay is set to , and the batch size is set to to optimize performance. For ResNet-50, the fine-tuning epochs are specifically set to , with distinct configurations for repeated and random augmentation, indicating a tailored approach to maximize the model’s efficacy on the ImageNet-1K challenge.
A.2.2 Semantic Segmentation Task
We followed the pipeline demonstrated by the iBot [Zhou et al., 2022] paper for fine-tuning the pre-trained model for semantic segmentation using the dataset [Zhou et al., 2017]. More specifically, we combined a pre-trained ViT-S/16 encoder with a UPerNet decoder [Xiao et al., 2018]. The ViT-S/16 encoder extracts detailed features from images, while the UPerNet decoder specializes in semantic segmentation, translating these features into precise pixel-level classifications. This process employed the AdamW optimizer and fine-tuned for iterations.
A.3 Projection Head
FOLK has three projection headers: one for frequency reconstruction (MFM Head, ) and two for the student () and teacher () header, see Figure 3. We used a single linear layer for the frequency reconstruction head, similar to that in the MFM method [Xie et al., 2023]. However, when it came to the distillation heads (student and teacher heads), we opted for a more sophisticated architecture akin to the DINO [Caron et al., 2021] head, albeit slightly modified. In FOLK, each head (student or teacher) comprised a 3-layer multi-layer perceptron (MLP) with a hidden dimensionality of 2048. All layers were followed by a GELU activation except the final layer. We refrained from applying batch normalization (BN), as ViT architectures, unlike standard CNNs, typically eschew BN by default. Rather than adhering to the 65536 output dimension in DINO, we followed the iBOT [Zhou et al., 2022] approach, adopting a dimensionality of 8192.
Appendix B Extra Experiments
B.1 Image Classification - CNN Base Modal
One limitation of many research studies, such as iBOT [Zhou et al., 2022] and AttMask [Kakogeorgiou et al., 2022], is that they are only compatible with a specific type of model architecture (e.g. ViTs), but not with others (e.g. CNNs). Our approach, like MFM [Xie et al., 2023], does not have this limitation and works with a wide range of encoder models.
Table 5 presents the ResNet-50 [He et al., 2016] (a CNN-based model) performance under the FOLK framework. The same FOLK pre-training strategies that have been applied to ViTs were seamlessly adopted here for CNNs. The only necessary modification to adopt CNNs is to alter the inputs to the heads: replacing patch tokens from ViTs with reshaped feature maps from CNNs, and replacing the token from ViTs with the average-pooled (and reshaped) feature map from CNNs. Our approach has demonstrated equal or superior performance to other methods with fewer epochs.
| Method | Ref | Epoch | Top-1 Acc |
| SimSiam | [Chen and He, 2021] | 400 | 79.1 |
| MoCo v2 | [Chen et al., 2020d] | 400 | 79.6 |
| SimCLR | [Chen et al., 2020b] | 800 | 79.9 |
[…]
| SwAV | [Caron et al., 2020] | 600 | 80.1 |
| MFM | [Xie et al., 2023] | 300 | 80.1 |
| MFM + Com/RCom | [Xie et al., 2023] | 300 | 80.1 |
| FOLK | Ours | 300 | 80.1 |
B.2 Few Shot Learning
In addition to the primary few-shot learning results discussed in Section 4.2.2, Table 6 presents an extensive evaluation of few-shot learning performance. Various pre-trained models were fine-tuned using only of the ImageNet-1K dataset over epochs. This setup facilitates a detailed comparison of each model’s ability to adapt to new data with minimal examples, highlighting a crucial aspect of model robustness and versatility. Furthermore, the evaluation considers three different settings for the base learning rate (LR) and warm-up periods, which are crucial hyperparameters in training deep learning models, especially under few-shot scenarios. The different configurations aim to assess each model’s robustness across varying learning rate adaptation conditions.
| Method | Base LR = 2e-4 | Base LR = 2e-4 | Base LR = 2e-3 | AVG |
| Warm Up = 0 | Warm Up = 100 | Warm Up = 5 | ||
[…]
| MFM | 26.9 | 31.6 | 6.3 | 21.6 |
[…]
| FOLK | 51.7 | 56.1 | 20.5 | 42.8 |
B.3 Ablation Study - Different Filters

### v2

@@ 4.2.2 Few Shot Learning
| BLR = 2e-4 | BLR = 2e-4 | BLR = 2e-3 | ||||
| Method | WUp = 0 | WUp = 100 | WUp = 5 | AVG | MAX | Epoch |
| iBOT | 64.0 | 71.1 | 2.0 | 45.7 | 71.1 | 800 |
| AttMask | 69.8 | 71.0 | 31.3 | 57.4 | 71.0 | 300 |
| MFM | 57.7 | 58.5 | 41.9 | 52.7 | 58.5 | 300 |
| MFM + Com/RCom | 66.3 | 63.9 | 59.9 | 63.4 | 66.3 | 300 |
| FOLK | 71.2 | 68.1 | 62.2 | 67.2 | 71.2 | 300 |
@@
| FOLK‡ | 77.4 | 74.9 | 66.4 | 72.9 | 77.4 | 800 |
The results in Table 2 highlight FOLK consistently outperforms methods like iBOT, AttMask, MFM, and MFM with Com/RCom filters across different base learning rates (BLR) and warm-up (WUp) settings. For instance, with a BLR of and no warm-up (WUp=0), FOLK achieves a top-1 accuracy of , surpassing iBOT’s , AttMask’s , MFM’s , and MFM with Com/RCom’s . When the pre-training epochs are increased to 800, FOLK’s performance further improves, reaching accuracy under the same BLR and WUp settings, and even achieving when fine-tuned for 1000 epochs (denoted as FOLK‡). The average accuracy across the three learning rate settings for FOLK with 800 pre-training epochs is , significantly higher than iBOT’s with the same number of pre-training epochs. Notably, FOLK maintains robust performance even when the learning rate is less optimal, such as achieving accuracy with a higher BLR of and a short warm-up period (WUp=5), where other methods like iBOT drop drastically to accuracy. These results demonstrate that FOLK not only excels in leveraging limited labeled data but also exhibits resilience to variations in fine-tuning hyperparameters, underscoring its effectiveness and adaptability compared to other state-of-the-art methods in few-shot learning scenarios.
5 Conclusion
Acknowledgments
This project was made possible, in part, by support from the National Science Foundation grant number OAC-2018627 and from the Ohio Supercomputer Center.
Appendix
The following appendices provide in-depth information to supplement our main findings and methodologies presented in the paper. These appendices aim to enrich the reader’s understanding by providing comprehensive insights and evidence supporting the robustness and versatility of our proposed method.
Appendix A: Implementation details like pre-training, fine-tuning process, and projection heads architecture.
Appendix B: Additional experimental results and analysis like semantic segmentation, few-shot learning, image classification for CNN-based model, and ablation study.
Appendix C: Distinctiveness of FOLK with other SSL based model.
Appendix D: Visual prediction results.
Appendix A Implementation Details
Throughout our experimental investigations, we adopted the methodologies prescribed by the iBOT (Zhou et al., 2022) and MFM (Xie et al., 2023) frameworks. In a multi-GPU training circumstance, the learning rate adjustment is vital for optimizing the model’s learning efficiency. The formulation used to compute the scaled learning rate is delineated below:
In this context, signifies the optimal or peak learning rate identified for the model’s training process. The term refers to the number of training examples processed simultaneously by each GPU. Lastly, denotes the total count of GPUs employed for parallel computation. The coefficient in the denominator is a normalization factor, ensuring the scaled learning rate maintains an appropriate magnitude relative to the hardware configuration. We used the PyTorch Library (Paszke et al., 2019) for our code development.
A.1 Pre-Train Stage
Our pre-training procedures largely align with those outlined in the BEiT (Bao et al., 2021) study, albeit with a few modifications. Specifically, our pre-training regime for the models incorporates simple yet effective data augmentation techniques, including random resized cropping to a resolution of pixels and image flipping.
We employ the AdamW optimizer (Loshchilov & Hutter, 2019), with a pre-training duration set to or epochs, a batch size of , per GPU, and a peak learning rate of . Additional parameters include a cosine decay learning rate schedule, warmup epochs, and a specific setting for optimizer momentum (, = , ) (Chen et al., 2020a) with a weight decay of . Also, we used a value of for gradient clipping to prevent the exploding gradient problem.
A.2 Fine-Tune Stage
For the fine-tuning stage, we tried to keep most of the configuration of our FOLK the same as MFM (Xie et al., 2023) for a fair comparison.
A.2.1 Classification Task
We ran epochs for fine-tuning the pre-trained model (i.e. ViT-S/16 or ViT-B/16) on ImageNet-1K for image classification, employing the optimizer across all configurations with a weight decay of and the optimizer momentum , = , . Moreover, the approach includes a cosine decay learning rate schedule (Li & Arora, 2020), with a layer-wise learning rate decay equal to (Bao et al., 2021; Clark et al., 2020). We also utilized advanced augmentation techniques such as Mixup (Zhang et al., 2018) and Cutmix (Yun et al., 2019), as well as label smoothing and random augmentation to further improve model robustness and generalization capability (Szegedy et al., 2016; Cubuk et al., 2020). The batch size is maintained at , with a peak learning rate set at .
In contrast, the fine-tuning settings for the ResNet-50 model (He et al., 2016) generally follow the configurations suggested by (Wightman et al., 2021), with modifications to adopt the AdamW optimizer as recommended by (Fang et al., 2023). This includes a binary cross-entropy loss function (Zhang & Sabuncu, 2018) and adjustments to the learning rate scheduler. The weight decay is set to , and the batch size is set to to optimize performance. For ResNet-50, the fine-tuning epochs are specifically set to , with distinct configurations for repeated and random augmentation, indicating a tailored approach to maximize the model’s efficacy on the ImageNet-1K challenge.
A.2.2 Semantic Segmentation Task
We followed the pipeline demonstrated by the iBot (Zhou et al., 2022) paper for fine-tuning the pre-trained model for semantic segmentation using the dataset (Zhou et al., 2017). More specifically, we combined a pre-trained ViT-S/16 encoder with a UPerNet decoder (Xiao et al., 2018). The ViT-S/16 encoder extracts detailed features from images, while the UPerNet decoder specializes in semantic segmentation, translating these features into precise pixel-level classifications. This process employed the AdamW optimizer and fine-tuned for iterations.
A.3 Projection Head
FOLK has three projection headers: one for frequency reconstruction (MFM Head, ) and two for the student () and teacher () header, see Figure 3. We used a single linear layer for the frequency reconstruction head, similar to that in the MFM method (Xie et al., 2023). However, when it came to the distillation heads (student and teacher heads), we comprised a 3-layer multi-layer perceptron (MLP) with a hidden dimensionality of 2048. All layers were followed by a GELU activation except the final layer. We refrained from applying batch normalization (BN), as ViT architectures, unlike standard CNNs, typically eschew BN by default. Also, the output dimension adopts a dimensionality of 8192.
A.4 Pseudocode of Informed Filters
Algorithm 1 presents the pseudocode for our proposed and filters (denoted as in Eq. 2 and Eq. 3). Here, is the set of predefined values ( in our experiments) from which a threshold is uniformly sampled to filter frequency components with the highest magnitudes. The parameter represents the probability of applying the filter, set to in our experiments; accordingly, the filter is applied with probability . Ablation studies evaluating different choices of and are provided in Appendix Sections B.5.1 and Section B.5.2, respectively.
Appendix B Extra Experiments
B.1 Semantic Segmentation
@@
Results of this fine-tuning effort for the semantic segmentation task on the ADE20K dataset are shown in Table 3. It presents a comparative analysis of different methodologies: Supervised Learning, iBOT, iBOT with Attention Mask (AttMask), MFM, and our FOLK model at different pre-training epochs ( and epochs). Notably, the FOLK model with epochs of pre-training achieves the highest mIoU at , slightly surpassing the iBOT’s mIoU. This indicates a successful adaptation of the FOLK methodology, showing not only an improvement over the MFM results but also demonstrating that extended pre-training can lead to marginal yet significant performance gains.
B.2 Image Classification - CNN Base Modal
One limitation of many research studies, such as iBOT (Zhou et al., 2022) and AttMask (Kakogeorgiou et al., 2022), is that they are only compatible with a specific type of model architecture (e.g. ViTs), but not with others (e.g. CNNs). Our approach, like MFM (Xie et al., 2023), does not have this limitation and works with a wide range of encoder models.
Table 4 presents the ResNet-50 (He et al., 2016) (a CNN-based model) performance under the FOLK framework. The same FOLK pre-training strategies that have been applied to ViTs were seamlessly adopted here for CNNs. The only necessary modification to adopt CNNs is to alter the inputs to the heads: replacing patch tokens from ViTs with reshaped feature maps from CNNs, and replacing the token from ViTs with the average-pooled (and reshaped) feature map from CNNs. Our approach has demonstrated equal or superior performance to other methods with fewer epochs.
| Method | Ref | Epoch | Top-1 Acc |
| SimSiam | (Chen & He, 2021) | 400 | 79.1 |
| MoCo v2 | (Chen et al., 2020d) | 400 | 79.6 |
| SimCLR | (Chen et al., 2020b) | 800 | 79.9 |
[…]
| SwAV | (Caron et al., 2020) | 600 | 80.1 |
| MFM | (Xie et al., 2023) | 300 | 80.1 |
| MFM + Com/RCom | (Xie et al., 2023) | 300 | 80.1 |
| FOLK | Ours | 300 | 80.1 |
In the experiments above, we observed that FOLK did not lead to significant gains over MFM when using CNN-based models in the standard many-shot (full-data) setting. However, upon further investigation, we discovered that FOLK offers considerable advantages over MFM in few-shot learning scenarios, even with CNN architectures. We conducted additional experiments using a ResNet-50 (He et al., 2016) model as the backbone for both FOLK and MFM, and evaluated their performance on ImageNet-1k with limited labeled data ( and subsets). The experimental setup mirrored that of our ViT-based models to ensure a fair comparison.
| Model | Ref | Epoch | 10% | 1% |
| MFM | (Xie et al., 2023) | 300 | 53.9 | 22.7 |
| MFM + Com/RCom | (Xie et al., 2023) | 300 | 58.6 | 31.0 |
| FOLK | Ours | 300 | 71.1 | 46.2 |
As shown in Table 5, FOLK significantly outperforms MFM in the few-shot setting when using a CNN backbone. Specifically, FOLK achieves a top-1 accuracy of with labeled data and with labeled data, compared to MFM’s and , respectively. These results highlight FOLK’s robustness and effectiveness in scenarios with limited labeled data, making it a superior choice for applications where labeled data is scarce.
The performance gains can be attributed to FOLK’s self-distillation mechanism, which allows the student model to learn from the teacher model’s representations of unmasked images. This mechanism provides additional guidance during pre-training, enabling the student model to better understand the underlying data distribution, even when the amount of labeled data for fine-tuning is limited. In contrast, MFM’s reliance solely on reconstructing masked frequencies without exposure to unmasked images during pre-training may limit its effectiveness in low-data regimes.
In the full-data setting, the performance gap between FOLK and MFM using CNNs is less pronounced. This observation aligns with prior findings that CNNs, due to their inherent inductive biases and localized receptive fields, may not benefit as much from global context as ViT architectures do. Nonetheless, the substantial improvements observed in few-shot settings demonstrate FOLK’s potential to enhance CNN-based models, especially when labeled data is limited.
These findings underscore the importance of evaluating self-supervised learning methods across different architectures and data availability scenarios. By effectively incorporating the original image information through self-distillation, FOLK enhances the pre-training process for CNNs, leading to better generalization in downstream tasks with scarce labeled data.
B.3 Robustness to Image Noise and Artifacts
While our FOLK framework leverages the Fourier transform to create frequency-masked augmentations, concerns may arise regarding the potential sensitivity of the Fourier transform to image artifacts and noise. It is well-known that the Fourier transform can amplify the impact of noise due to its global nature (Oppenheim & Lim, 1981). However, our method uses the Fourier transform primarily to generate diverse and challenging masked inputs, and the actual model training occurs in the spatial domain.
To assess the robustness of our model’s learned representations of image noise and artifacts, we conducted an experiment evaluating the impact of common types of image degradation on the model’s performance in reconstructing masked frequencies. We randomly selected 100 images from the ImageNet-1k validation set and applied following types of degradation to each image:
Salt-and-Pepper Noise: Randomly replaces pixels with maximum or minimum intensity values, simulating impulsive noise.
Gaussian Noise: Adds zero-mean Gaussian noise with a standard deviation of to simulate sensor noise.
Brightness Variation: Adjust the brightness by a random factor within to simulate different lighting conditions.
Contrast Variation: Adjust the contrast by a random factor within to mimic variations in image quality.
For each original and degraded image, we applied the or filters described in Section 3.2.1 to create frequency-masked versions. We then used our pre-trained FOLK model to predict the masked frequencies and calculated the MFM loss as defined in Equation 3.
| Noise Type | Without Noise | Salt-and-Pepper | Gaussian | Brightness | Contrast |
| MFM Loss |
Table 6 presents the average MFM loss values for the original images and images subjected to various types of degradations. The results indicate that these degradations cause only a slight increase in the MFM loss—approximately to higher than the loss for the original images. This marginal difference suggests that common image degradations do not significantly compromise our model’s ability to reconstruct masked frequencies.
Several factors contribute to the robustness of our model to image artifacts and noise:
Spatial Domain Training: Although we use the Fourier transform for masking, the input to the model remains in the spatial domain. This approach mitigates the potential amplification of noise in the frequency domain.
Randomized Masking Parameters: The use of randomly sampled thresholds for filter creation (as discussed in Section 3.2.1) exposes the model to a wide range of frequency masking scenarios. This diversity encourages the model to focus on significant frequencies while reducing sensitivity to noise.
Data Augmentation: Our training pipeline includes standard data augmentations such as random cropping and flipping. These augmentations help the model learn robust features that generalize well to variations in the input data.
The minimal impact of noise on the MFM loss indicates that our model effectively learns to reconstruct masked frequencies even when the input images contain artifacts. This robustness is crucial for practical applications where images may be degraded due to sensor imperfections or environmental conditions.
B.4 Few Shot Learning
In addition to the primary few-shot learning results discussed in Section 4.2.2, Table 7 presents an extended evaluation of few-shot learning performance using a smaller set of labeled data. Various pre-trained models were fine-tuned using only of the ImageNet-1K dataset over epochs. This setup facilitates a detailed comparison of each model’s ability to adapt to new data with minimal examples, highlighting a crucial aspect of model robustness and versatility. Furthermore, the evaluation considers three different settings for the base learning rate (LR) and warm-up periods, which are crucial hyperparameters in training deep learning models, especially under few-shot scenarios. The different configurations aim to assess each model’s robustness across varying learning rate adaptation conditions.
| BLR = 2e-4 | BLR = 2e-4 | BLR = 2e-3 | ||||
| Method | WUp = 0 | WUp = 100 | WUp = 5 | AVG | MAX | Epoch |
| iBOT | 33.2 | 59.0 | 1.4 | 31.2 | 59.0 | 800 |
[…]
| MFM | 26.9 | 31.6 | 6.3 | 21.6 | 31.6 | 300 |
| MFM + Com/RCom | 42.5 | 44.5 | 10.7 | 32.6 | 44.5 | 300 |
[…]
| FOLK | 58.9 | 64.2 | 29.3 | 50.8 | 64.2 | 800 |
B.5 Ablation Study
B.5.1 Rationale Behind Threshold Selection
Our primary goal in selecting multiple threshold values was to introduce diversity in the frequency-masked input images during pre-training. By randomly sampling thresholds from a set (e.g., [0.005, 0.01, 0.05]), we ensure that the model is exposed to a variety of masking levels, which helps it learn more robust and generalized representations. The thresholds correspond to the proportion of frequency components retained or masked; for instance, a threshold of 0.005 means we retain (or mask) the top of the highest magnitude frequencies. By varying this threshold, we create different levels of difficulty for the reconstruction task, encouraging the model to learn both coarse and fine-grained features.
| Threshold(s) | CIFAR-10 | CIFAR-100 | ImageNet-1k |
| .05 | 98.8 | 90.2 | 81.1 |
| .01 | 99.0 | 90.7 | 81.4 |
| .005 | 99.0 | 90.6 | 81.3 |
| .005,.01,.05 | 99.1 | 90.9 | 81.6 |
| .002,.01,.07 | 99.1 | 90.9 | 81.6 |
While the initial selection of these thresholds might seem arbitrary, our experiments indicate that the exact values are not highly sensitive parameters. To demonstrate this, we conducted additional experiments using different sets of thresholds and evaluated their impact on downstream task performance. We compared the original set of thresholds [0.005, 0.01, 0.05] with an alternative set [0.002, 0.01, 0.07]. The results, presented in Table 8, show that both sets yield similar performance across various datasets (CIFAR-10, CIFAR-100, and ImageNet-1k), and both outperform using a single threshold value.
These results suggest that the method’s effectiveness is robust to the specific choice of thresholds as long as they are sufficiently separated to produce diverse frequency-masked images. The performance gains from using multiple thresholds indicate that the model benefits from the increased variability in the masking patterns. In practice, one can select a set of thresholds that produce visually distinct frequency-masked images without extensive tuning. We found that thresholds leading to the retention or masking of approximately to of the highest magnitude frequencies work well. Visual inspection of the resulting images can help in determining appropriate thresholds, as shown in Figures 13, 14, and 15 in the Appendix. By observing the images produced with different thresholds, practitioners can ensure that the selected thresholds provide a good balance between retaining important image structures and introducing sufficient masking to challenge the model during pre-training. The choice of multiple thresholds enhances the diversity of the training data without requiring extensive hyperparameter tuning. Our empirical results demonstrate that as long as the thresholds are spread out to produce varying levels of frequency masking, the model performance remains robust. This approach avoids the need for fine-grained tuning of threshold values and simplifies the practical application of our method.
In summary, while we initially selected the thresholds with the intention of creating diverse and challenging pre-training tasks, our experiments indicate that the exact values are not critical. The key is to have multiple thresholds that lead to different masking levels, and our findings show that the method is effective across various datasets with minimal sensitivity to these values.
B.5.2 Effect of Filter Selection Probability
To assess the impact of different selection probabilities for the and filters in our FOLK framework, we conducted an ablation study by varying the probability of selecting each filter during pre-training. While our default setting randomly selects between the two filters with equal probability , this experiment aims to determine whether this choice is optimal and how sensitive our method is to this hyperparameter.
| Com Probability | RCom Probability | Accuracy |
| 0.1 | 0.9 | 81.3 |
| 0.3 | 0.7 | 81.4 |
| 0.5 | 0.5 | 81.6 |
| 0.7 | 0.3 | 81.4 |
| 0.9 | 0.1 | 81.4 |
The top-1 accuracy results on the ImageNet-1k validation set are presented in Table 9 We observe that selecting the filters with equal probability () yields the highest accuracy of . Slight deviations from this equal probability result in marginal decreases in performance, with accuracies ranging from to . The performance differences are within of the highest accuracy, indicating that our method is not highly sensitive to the exact selection probability of the filters.
These results suggest that while an equal selection probability slightly outperforms other settings, the FOLK framework is robust to variations in the selection probability of the and filters. The performance does not significantly degrade when the selection probabilities are skewed towards one filter over the other. This indicates that both filters contribute positively to the learning process, and the model benefits from the diversity introduced by both types of frequency masking.
B.5.3 Weight Optimization
@@
B.5.4 Different Filters